# A is for Absorption: Studying Feature Splitting and Absorption in Sparse Autoencoders

### v2

@@ 1 Introduction
One recent promising approach in this field is the use of Sparse Autoencoders (SAEs), which have shown potential in decomposing the dense, polysemantic activations of LLMs into more “interpretable” latent features (Huben et al., 2024; Bricken et al., 2023) using sparse dictionary learning (Olshausen & Field, 1997). SAE neurons (hereafter called “latents”)  [^1] We use latents to prevent overloading the term feature, which we reserve for human-interpretable concepts the SAE may capture. This breaks from earlier usage which used feature for both (Elhage et al., 2022), but aligns with the terminology in (Lieberum et al., 2024) and makes the distinction more clear. are said to be interpretable if they appear to detect some property of the input (which we refer to as a “feature”) and they classify that feature with high precision / recall (Bricken et al., 2023).
The ability to decompose LLM activations into monosemantic latents presents an opportunity to not only interpret but also modify internal model representations for various downstream tasks (Marks et al., 2024). However, despite these theoretical advantages, most demonstrations of SAEs for interpretability have been limited to toy models (Sharkey et al., 2022) or concepts that are difficult to validate empirically (Huben et al., 2024). Bricken et al. (2023) performs some case-studies on SAE latents in a one-layer transformer model, but assumes that gaps in recall are minor and due to imperfections in the model or data and does not investigate further.
@@ 6 Discussion
We hope as well this investigation may lead to research into solutions, particularly those involving Meta-SAEs (Bussmann et al., 2024), to solve or mitigate feature absorption. Anoter possible solution may be attribution dictionary learning (Olah et al., 2024).

### v3

@@ 1 Introduction
One recent promising approach in this field is the use of Sparse Autoencoders (SAEs), which have shown potential in decomposing the dense, polysemantic activations of LLMs into more “interpretable” latent features (Huben et al., 2024; Bricken et al., 2023) using sparse dictionary learning (Olshausen & Field, 1997). SAE neurons (hereafter called “latents”)  [^1] We use latents to prevent overloading the term feature, which we reserve for human-interpretable concepts the SAE may capture. This breaks from earlier usage which used feature for both (Elhage et al., 2022), but aligns with the terminology in (Lieberum et al., 2024) and makes the distinction more clear. are said to be interpretable if they appear to detect some property of the input (which we refer to as a “feature”) and classify that feature with high precision / recall (Bricken et al., 2023).
However, despite these theoretical advantages, most demonstrations of SAEs for interpretability have been limited to toy models (Sharkey et al., 2022) or concepts that are difficult to validate empirically (Huben et al., 2024). Subsequent analysis measures recall of SAE latents with dataset filtering and LLM-based analysis and notes that cooperation between many latents may be required to account for all instances of a concept (Olah et al., 2024b). We build on this work by evaluating precision / recall on a large number of Gemma Scope SAEs, and offer a partial explanation for lower-than-expected recall of SAE latents in the form of “feature absorption”.
@@ 5.1 Applications of Probes and SAEs for Model Interpretability
Other work has noted poor recall / precision of SAE latents compared to known proxies (Olah et al., 2024b; Kissane et al., 2024; Templeton et al., 2024). We build on this work by evaluating a large number of Gemma Scope SAEs to demonstrate how precision / recall is mediated by sparsity, and also offer a possible explanation of low recall due to feature absorption.
@@ 6 Discussion
We hope as well this investigation may lead to research into solutions, particularly those involving Meta-SAEs (Bussmann et al., 2024), to solve or mitigate feature absorption. Anoter possible solution may be attribution dictionary learning (Olah et al., 2024a).
@@ A.1 Glossary of Terms
Circuit:
In the context of neural network interpretability, a circuit refers to a subgraph of neurons or features within a neural network that work together to perform a specific function or computation. The study of circuits aims to understand how different components of a neural network interact to process information and produce outputs.